OpenStudy (anonymous):

Evaluate each polynomial for x = -1 and y = 3. x + 3y

OpenStudy (hoblos):

replace x by -1 and y by 3 x + 3y = (-1) +3(3) = -1+9 = 8

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

-1+ 3(3)

OpenStudy (hoblos):

my pleasure :)